# The LCLStream Ecosystem for Multi-Institutional Dataset Exploration

### v1

@@ Abstract.
conference: XLOOP 2025; November, 2025; St. Louis, MO [^1] The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan).
@@ 1. Introduction
Autonomous experiment steering is already needed now to quickly adapt to changing experimental conditions that guide the instrument to optimal operating regimes. For example, machine-driven, Bayesian optimization of mechanical alignment settings for the incoming X-ray waveguides achieves 5x faster time-to-calibration than manual, human-driven optimization.(Roussel et al., 2024)
However, coupling High-Performance Computing (HPC) to external, on-line data sources requires the convergence of several new capabilities: data ontologies(Rajamohan et al., 2025) for the naming, processing, and storage of results at each level (e.g. raw events, X-ray crystal images, detected scattering peak positions, or electron detection times), experimental analysis frameworks for interactively guiding (CPU/GPU-intensive) data exploration, cross-facility coordination on co-scheduling of experiments and analysis, web-accessible applications programming interfaces (API-s) for data producers and HPC jobs, robust software for high-bandwidth data transmission, and a programming language coordinating interlinked, multi-participant activities.
@@ 1.1. Data Streaming Framework
Experimental data at LCLS is accessed using the psana1 and psana2 libraries (for LCLS and LCLS-II respectively) (Damiani et al., 2016; Thayer et al., 2024). These libraries read and assemble events from a running or archived experiment in a highly parallelized fashion. These libraries are scalable, performant and flexible, but must be used on site for administrative and technical reasons. The services that provide data access run user code and access databases (containing information needed to pre-process events). These actions can only be performed by authenticated users from within the LCLS facility.
The streaming framework described in this work opens up these data by providing local and remote distributed access to the psana1 and psana2 data outputs. Preliminary timing results show data arrival at an HPC job running on Oak Ridge’s National Center for Computational Sciences just seconds after collection at SLAC’s detectors in Menlo Park. The net result is increased data availability for the growing community of experts whose work focuses on experimental steering, time-sensitive processing, and online analysis.(Rakitin et al., 2020; Brim et al., 2024; Mariani et al., 2016)
Figure 1. Data streaming process diagram. Blue arrows show control paths, and black arrows show data flow. Dotted paths are for returned results. Event assembly and data formatting is performed by the psana framework inside S3DF (left). The LCLStream API can start network buffers and MPI jobs on S3DF to format and send experimental data. External users can pair an LCLStream API call with jobs on other HPC clusters (right).
Fig. 1 shows the core components of our data streaming framework. The LCLSstream API Server (in blue at the center of figure 1) allows beamline or external users to request a dataset from a specific experiment using a REST API with JSON-formatted queries. The user performing the request and the server recognize each other through a certificate-driven mutual authentication process. LCLStream-API creates a JobID and launches the LCLStreamer and NNG-Stream components of the pipeline on LCLS’s S3DF cluster (on the left in the same figure).
@@ 2. Scientific Applications
Given this historical experience, the OM package for crystallographic data analysis has built a specification for a data pre-processing pipeline.(Mariani et al., 2016) In order to adapt this idea to LCLStreamer, we considered the needs of several new and existing data analysis projects:
@@
Specifically, we looked for patterns in the data reduction steps and the input data formats used by each application. We also considered the needs of several developing projects which could make use of this data processing pipeline, including: online calibration estimation methods, high-rate image transmission from the XPP hutch to GPU nodes for image processing, and custom setups for CCTBX users who want more control over their data formatting (e.g. transformation to NeXus data format).
@@
At the beginning of our work, all of the above applications were able to run locally at the S3DF, by reading xtc or xtc2-formatted event data using LCLS’s psana package.(Damiani et al., 2016; Thayer et al., 2024) These event data files are output by each experimental run, can be generated at rates up to gigabytes per second (which will become terabytes in the future), and are effectively stored/loaded/streamed in parallel by LCLS’s high-throughput hardware setup. As the project progressed, we merged the common steps in these applications, and abstracted the differences into configurable options. Most variations can now be handled by adding new input detectors and data reduction functions, rather than rebuilding the entire data processing application.
@@ 2.1. Masked X-ray Image Autoencoder (MAXIE) and PeakNet
Cong and Chen describe two AI models capable of reconstructing and interpreting X-ray images.(Wang et al., 2025; Chen et al., 2025) The team has developed the Masked Autoencoder for X-ray Image Encoding (MAXIE), supporting model architectures ranging from hundreds of millions to billions of parameters. MAXIE is trained on approximately 286 terabytes of X-ray diffraction images from a variety of different detectors. Training epochs requiring hours to days depending on available compute resources. The implementation supports multiple parallelization strategies within a unified training framework, including single GPU, multi-GPU, and multi-node configurations using both Distributed Data Parallel (DDP) and Fully Sharded Data Parallel (FSDP) approaches (including sharded and full checkpoints), with optimizations including shared memory utilization and job scheduler integration for fault-tolerant execution.
The goal for this application is to run unsupervised training of this model on the full 286-terabyte dataset, potentially unlocking representations learned from the complete diversity of crystallographic data. Current production results come from a complementary supervised learning approach operating on a significantly smaller, algorithmically labeled dataset. Because both approaches leverage the same unified training infrastructure, scaling up to larger datasets is a matter of compute availability and network data accessibility. Architectural consistency in the AI/ML framework provides operational benefits including shared fault-tolerance mechanisms, unified checkpoint management, and the ability to scale supervised training to larger datasets as they become available. Preliminary results suggest that we need robust, flexible methods like LCLStreamer for pre-processing our data as we work to improve its quality.
@@ 2.2. TMO-prefex
The LCLS-II upgrade for the Time-resolved atomic, Molecular and Optical Science (TMO) beamline is designed to operate with X-ray laser shot repetition rates up to 1 MHz, collecting electron time of flight data with sub-femtosecond resolution after each shot.(Walter et al., 2021)
In theory, the increase in shot repetition rate from the previous 120 Hz allows experiments to complete 8333 times faster. However, the current state of the art for data analysis is running offline, out of the main experimental path. When done this way, it is not possible to determine whether each experimental step has been correctly configured and informative data has been gathered fast enough to keep up with the speed of data generation. In order to actually accelerate experimental the data collection phase, we need a fast data analysis pipeline. Speeding up the iteration time needed for completing each experimental step is key to realizing 1000x productivity increases. It will ultimately require data analysis capable of reconfiguring the experimental parameters in an automated way.(Roussel et al., 2024)
@@
Figure 2. TMO time of flight (ToF) detector configuration for detecting time and angular distribution of emitted electrons (reproduced from Ref. (Gouin-Ferland et al., 2022)). ToF spectrometer signals are processed by analog electronics before being digitized. After event detection, the central FPGA (circled with a red dotted line), forwards event features from all 8 peripheral FPGA-s on to the S3DF data processing pipeline.
@@ 2.3. CrystFEL
CrystFEL (White et al., 2012) is a collection of programs for processing data collected using the Serial Femtosecond Crystallography technique. It includes programs performing all the steps needed to turn a set of diffraction images in a list of structure factors. This includes utilities to index and integrate diffraction patterns, and to merge measured intensities. CrystFEL can operate both on images stored in HDF5 format(Group, 2025) in local files, or streamed via a network socket in the same format. Typically a data packet includes a diffraction image, and supplemental information needed for its interpretation (an estimation of the distance between the detector and the interaction point of the experiment, approximate beam energy, optical laser status, etc). Some additional pertinent experimental information (detector geometry, space group symmetry, etc.) is instead provided via program configuration files or command-line parameters. A graphical front-end to control the processing is available.
During a beamtime at the Macromolecular Femtosecond Crystallography (MFX) endstation at LCLS, the LCLSstreamer component was used to send a data from one of the experimental runs over to the testbed HPC cluster at the Oak Ridge Leadership Computing Facility, where CrystFEL had been setup to receive and process the transferred data. The latency between data collection and processing with CrystFEL has been shown to be within the range of 15-25 seconds, allowing useful feedback information within a time range sufficient for researchers to manually steer the experiment in the desired direction.
Distributing this processing pipeline is somewhat simplified by the program’s ability to read source data from the network. However, it is still challenging to arrange sending of the extra metadata while starting and stopping the program on an HPC center. In addition, CrystFEL stores its intermediate and final outputs as files on the computer where it is run. Hence there is a need for a return channel to send results back to the user’s file location.
The state of the art for crystallography workflows is still to run the graphical user interface on the HPC compute login node, and attach to it using a remote desktop.(Blaschke et al., 2024). The workflow aspect was handled by setting up the GUI to launch SLURM scripts.
Although CrystFEL can also start SLURM jobs and run containerized in HPC environments, we envision a fully remote task submission system to support experiments with CrystFEL and other cystallography data processing packages like CCTBX (Brewster et al., 2025) in the near future. Psi-K (described below) both has its own API, and can make use of other job APIs like NERSC SFAPI,(Enders et al., 2020) S3M,(Skluzacek et al., 2025) and Globus.(Ananthakrishnan et al., 2024) to support task submission system on the remote machine. Asynchronous, event-driven callbacks will keep the front-end informed of completion of each step of the processing pipeline. Automation and user interaction in the controller with then schedule successive operations. At each step, output files can be pulled (or streamed) back to the controller for display to the local user.
@@ 3. Software Infrastructure
LCLStreamer was designed based on patterns proven in the OM X-ray analysis software suite.(Mariani et al., 2016) This package has been developed over several years and has a strong user base within the X-ray science community. Its interface is designed around a top-level specification file for the data collection, reduction, and analysis pipeline. Users are able to achieve flexibility by plugging in different producers, data processing steps, and consumers within that pipeline.
@@
The state sequence and the format of the data structures used by Psi-k benefit strongly from specifications developed by the ExaWorks Psi/J project.(Hategan-Marandiuc et al., 2023) Psi-k’s design diverged, however, in order to provide web interoperability. For example, state changes are stored in a status file, and can also trigger webhooks. Stdout and stderr are captured in a logs folder, and numbered sequentially for each re-run of the job. These file layouts form the basis for exposing jobs via an API (Psik-API). Mapping job create-read-update-destroy operations to the file hierarchy provides a straightforward REST-HTTPS interface to Psi-k.
@@ 3.1. Fast Data Reduction with LCLStreamer
The LCLSstreamer application is made up of several parts:
@@
The standard pipeline batches together the results of processing several consecutive events. This accomplishes the same kind of batching one sees in a pytorch DataLoader, but with more configurability in the pipeline steps. At that point, the accumulated data is returned in bulk. The data still has the format of a dictionary of Numpy array, with each key representing a data entry, and the corresponding value storing the accumulated data. The data is then serialized into a binary form.
@@
The lclstreamer section of the configuration file determines which implementation of each component it should use. Each component entry in this section identifies a python class that implements the operations required by the component’s interface. For example:
lclstreamer:
  event_source: Psana1EventSource
  processing_pipeline:
    BatchProcessingPipeline
  data_serializer: Hdf5Serializer
  data_handlers:
    - BinaryFileWritingDataHandler
    - BinaryDataStreamingDataHandler
With these configuration options, LCLStreamer reads psana1 data (Psana1EventSource), batches the data
(BatchProcessingPipeline), serializes the data in a binary blob with the internal structure of an HDF5 file (Hdf5Serializer) and finally hands the binary blob to two data handlers: one that saves it as a file (BinaryFileWritingDataHandler) and one that streams it through a network socket (BinaryDataStreamingDataHandler).
Configuration options can be provided for each of the Python classes that implement the LCLStreamer components. For example, the configuration parameters for the Hdf5Serializer class, which implements the data serializer component, are defined by the Hdf5Serializer entry in the data_serializer section:
@@
Finally, the data_sources section of the configuration file defines the data that LCLStreamer extracts from every data event it processes. If a piece of information is part of the data event, but not included in the data_sources section, LCLStreamer will ignore it (filtering at read time). The data_sources section of the configuration file consists of a dictionary of data sources. Each entry has a key, which acts as a name that identifies the extracted data throughout the whole LCLStreamer data workflow, and a value, which is itelf a dictionary. This inner dictionary defines the nature of the data source (via the mandatory type entry) and any other parameters needed to configure it. The type of a data source is the name of the Python class that implements it.
For example:
@@
one called timestamp and one called detector_data. The timestamp data class is of type Psana1Timestamp. Inside LCLStreamer, the Psana1Timestamp class will be called to read the associated timestamp data for each event. The detector_data class is instead of type Psana1AreaDetector. The two configuration parameters psana_name and calibration are passed to the Python class Psana1AreaDetector that defines how this type of data is retrieved.
LCLStreamer currently supports psana (Damiani et al., 2016) and psana2 (Thayer et al., 2024) as data sources, with a wide array of detectors and instrument readouts. It allows data to be serialized in customizable formats like HDF5 and can write them to files and/or send them via ZMQ or NNG sockets.

### v2

@@ Abstract.
Streaming, X-Ray Detectors, Data Analysis, Integrated Research Infrastructure
journalyear: 2026copyright: ccconference: SCA/HPCAsia 2026 Workshops: Supercomputing Asia and International Conference on High Performance Computing in Asia Pacific Region Workshops; January 26–29, 2026; Osaka, Japanbooktitle: SCA/HPCAsia 2026 Workshops: Supercomputing Asia and International Conference on High Performance Computing in Asia Pacific Region Workshops (SCA/HPCAsiaWS 2026), January 26–29, 2026, Osaka, Japandoi: 10.1145/3784828.3786261isbn: 979-8-4007-2328-5/2026/01ccs: Information systems Personalizationccs: Information systems Data encoding and canonicalizationccs: Information systems RESTful web servicesccs: Information systems Extraction, transformation and loading
@@ 1. Introduction
Autonomous experiment steering is already needed now to quickly adapt to changing experimental conditions that guide the instrument to optimal operating regimes. For example, machine-driven, Bayesian optimization of mechanical alignment settings for the incoming X-ray waveguides achieves 5x faster time-to-calibration than manual, human-driven optimization.(Roussel and others, 2024)
However, coupling High-Performance Computing (HPC) to external, on-line data sources requires the convergence of several new capabilities: data ontologies(Rajamohan and others, 2025) for the naming, processing, and storage of results at each level (e.g. raw events, X-ray crystal images, detected scattering peak positions, or electron detection times), experimental analysis frameworks for interactively guiding (CPU/GPU-intensive) data exploration, cross-facility coordination on co-scheduling of experiments and analysis, web-accessible applications programming interfaces (API-s) for data producers and HPC jobs, robust software for high-bandwidth data transmission, and a programming language coordinating interlinked, multi-participant activities.
@@ 1.1. Data Streaming Framework
Experimental data at the Linac Coherent Light Sources (LCLS) is accessed using the psana1 and psana2 libraries (for LCLS and LCLS-II respectively).(Damiani and others, 2016; Thayer and others, 2024) These libraries read and assemble events from a running or archived experiment. These libraries are flexible, performant, and scalable across nodes using MPI, but must be used on site for administrative and technical reasons. The services that provide data access run user code and access databases (containing information needed to pre-process events). These actions can only be performed by authenticated users from within the LCLS facility.
The streaming framework described in this work opens up these data by providing local and remote distributed access to the psana1 and psana2 data outputs. Preliminary timing results show data arrival at an HPC job running on Oak Ridge’s National Center for Computational Sciences just seconds after collection at SLAC’s detectors in Menlo Park. The net result is increased data availability for the growing community of experts whose work focuses on experimental steering, time-sensitive processing, and online analysis.(Rakitin and others, 2020; Brim et al., 2024; Mariani and others, 2016)
Figure 1. Data streaming process diagram. Blue arrows show control paths, and black arrows show data flow. Dotted paths are for returned results. Event assembly and data formatting is performed by the psana framework on a compute job inside S3DF (left). The LCLStream API can start network buffers (on a S3DF data transfer node) and compute jobs on S3DF to format and send experimental data. External users can pair an LCLStream API call with jobs on other HPC clusters (right).
Fig. 1 shows the core components of our data streaming framework. The LCLStream API Server (in blue at the center of figure 1) allows beamline or external users to request a dataset from a specific experiment using a REST API with Javascript object notation (JSON)-formatted queries. The user performing the request and the server recognize each other through a certificate-driven mutual authentication process. LCLStream-API creates a JobID and launches a pair of processes on S3DF (the SLAC Shared Science Data Facility). First, it starts an NNG-Stream process a data transfer node with a new pair of recieve/send ports for this stream. Second, it starts LCLStreamer as an MPI job over SLURM on the S3DF cluster (on the left in the same figure). The receive URI is returned to the client. All compute processes can make independent connections to that address.
@@ 2. Scientific Applications
Given this historical experience, the OM (OnDA Monitor) package for crystallographic data analysis has built a specification for a data pre-processing pipeline.(Mariani and others, 2016) In order to adapt this idea to LCLStreamer, we considered the needs of several new and existing data analysis projects:
@@
Specifically, we looked for patterns in the data reduction steps and the input data formats used by each application. We also considered the needs of several developing projects which could make use of this data processing pipeline, including: online calibration estimation methods, high-rate image transmission from the X-Ray pump-probe (XPP) hutch to GPU nodes for image processing, and custom setups for CCTBX users who want more control over their data formatting (e.g. transformation to NeXus data format).
@@
At the beginning of our work, all of the above applications were able to run locally at the S3DF, by reading xtc or xtc2-formatted event data using LCLS’s psana package.(Damiani and others, 2016; Thayer and others, 2024) These event data files are output by each experimental run, can be generated at rates up to gigabytes per second (which will become terabytes in the future), and are effectively stored/loaded/streamed in parallel by LCLS’s high-throughput hardware setup. As the project progressed, we merged the common steps in these applications, and abstracted the differences into configurable options. Most variations can now be handled by adding new input detectors and data reduction functions, rather than rebuilding the entire data processing application.
@@ 2.1. Masked X-ray Image Autoencoder (MAXIE) and PeakNet
Cong and Chen describe two AI models capable of reconstructing and interpreting X-ray images.(Wang et al., 2025; Chen et al., 2025) The team has developed the Masked Autoencoder for X-ray Image Encoding (MAXIE), supporting model architectures ranging from hundreds of millions to billions of parameters. MAXIE is trained on approximately 286 terabytes of X-ray diffraction images from a variety of different detectors. Training epochs require hours to days depending on available compute resources, motivating its checkpointing and fault tolerance features. The implementation supports multiple parallelization strategies within a unified training framework, including single GPU, multi-GPU, and multi-node configurations using both Distributed Data Parallel (DDP) and Fully Sharded Data Parallel (FSDP) approaches (including sharded and full checkpoints), with optimizations including shared memory utilization and job scheduler integration for fault-tolerant execution.
The goal for this application is to run unsupervised training of this model on the full 286-terabyte dataset, potentially unlocking representations learned from the complete diversity of crystallographic data. Current production results come from a complementary supervised learning approach operating on a significantly smaller, algorithmically labeled dataset. Because both approaches leverage the same unified training infrastructure, scaling up to larger datasets is a matter of compute availability and network data accessibility. Architectural consistency in the AI/ML framework provides operational benefits including shared fault-tolerance mechanisms, unified checkpoint management, and the ability to scale supervised training to larger datasets as they become available.
@@ 2.2. TMO-prefex
The LCLS-II upgrade for the Time-resolved atomic, Molecular and Optical Science (TMO) beamline is designed to operate with X-ray laser shot repetition rates up to 1 MHz, collecting electron time of flight data with sub-femtosecond resolution after each shot.(Walter and others, 2021)
In theory, the increase in shot repetition rate from the previous 120 Hz allows experiments to complete 8333 times faster. However, the current state of the art for data analysis is running offline, out of the main experimental path. When done this way, it is not possible to determine whether each experimental step has been correctly configured and informative data has been gathered fast enough to keep up with the speed of data generation. In order to actually accelerate the experimental data collection phase, we need a fast data analysis pipeline. Speeding up the iteration time needed for completing each experimental step is key to realizing 1000x productivity increases. It will ultimately require data analysis capable of reconfiguring the experimental parameters in an automated way.(Roussel and others, 2024)
@@
Figure 2. TMO time of flight (ToF) detector configuration for detecting time and angular distribution of emitted electrons (reproduced from Ref. (Gouin-Ferland et al., 2022)). ToF spectrometer signals are processed by analog electronics before being digitized. After event detection, the central FPGA (circled with a red dotted line), forwards event features from all 8 peripheral FPGA-s on to the S3DF data processing pipeline. Instead of a custom code juggling event extraction with LCLS-II’s internal data path, LCLStream simplifies data extraction from this source down to a list of detector names and compression options.
@@ 2.3. CrystFEL
CrystFEL (White et al., 2012) is a collection of programs for processing data collected using the Serial Femtosecond Crystallography technique. It includes programs performing all the steps needed to turn a set of diffraction images in a list of structure factors. This includes utilities to index and integrate diffraction patterns, and to merge measured intensities.
Live data analysis as the user is taking data at a beamline is a killer feature because it allows the user to maximize the scientific output from their beamtime.(Brewster et al., 2025) For example, determining whether detectors are aligned, diffraction spot intensities cover a large enough dynamic range, or whether the crystal itself contains growth defects strongly impact the timeline of an experiment.
In support of this, CrystFEL can operate both on images stored in HDF5 format(Group, 2025) in local files, or streamed via a network socket in the same format. Typically a data packet includes a diffraction image, and supplemental information needed for its interpretation (an estimation of the distance between the detector and the interaction point of the experiment, approximate beam energy, optical laser status, etc). Some additional pertinent experimental information (detector geometry, space group symmetry, etc.) is instead provided via program configuration files or command-line parameters. A graphical front-end to control the processing is available.
Previous works showing live, HPC-assisted data analysis mirrored raw (xtc) data files to the remote facility and then ran the full psana-based software load process.(Blaschke and others, 2024) In the present work, we aim to do the extraction and compression in parallel at the source. Separating the data pipeline this way is somewhat simplified by the program’s ability to read source data from the network. However, it is still challenging to arrange sending of the extra metadata while starting and stopping the program on an HPC center. If successful, the receiving analysis software will be less complicated and the entire pipeline will run with lower latency and higher throughput.
@@ 3. Software Infrastructure
LCLStreamer was designed based on patterns proven in the OM X-ray analysis software suite.(Mariani and others, 2016) This package has been developed over several years and has a strong user base within the X-ray science community. Its interface is designed around a top-level specification file for the data collection, reduction, and analysis pipeline. Users are able to achieve flexibility by plugging in different producers, data processing steps, and consumers within that pipeline.
@@
The state sequence and the format of the data structures used by Psi-k benefit strongly from specifications developed by the ExaWorks PSI/J project.(Hategan-Marandiuc and others, 2023) Psi-k’s design diverged, however, in order to provide web interoperability. For example, Psi-k does not run job status checks by polling SLURM commands. State changes are stored in a status file, and can also trigger webhooks. Stdout and stderr are captured in a logs folder, and numbered sequentially for each re-run of the job. These file layouts form the basis for exposing jobs via an API (Psik-API). Mapping job create-read-update-destroy operations to the file hierarchy provides a straightforward REST-HTTPS interface to Psi-k.
@@ 3.1. Fast Data Reduction with LCLStreamer
The LCLStreamer application is made up of several parts:
@@
The standard pipeline batches together the results of processing several consecutive events. This accomplishes the same kind of batching one sees in a pytorch DataLoader, a key enabler of performance and interoperability. At that point, the accumulated data is returned in bulk. The data still has the format of a dictionary of Numpy array, with each key representing a data entry, and the corresponding value storing the accumulated data. The data is then serialized into a binary form.
@@
Since there are multiple choices for event sources, processing pipelines, data serializers, etc., each section of the configuration file is marked with a type to select which implementation class it will use.
Next to the type selection, additional configuration options can also be provided to each of these implementation classes. For example, the HDF5Serializer class implements the data serializer interface. It serializes its input data into a binary string with the internal structure of an HDF5 file. Its configuration options can defined be when the data_serializer section selects the HDF5Serializer type.
@@
Finally, the data_sources section of the configuration file defines the data that LCLStreamer extracts from every data event it processes. If a piece of information is part of the data event, but not included in the data_sources section, LCLStreamer will ignore it (filtering at read time). The data_sources section of the configuration file consists of a dictionary of data sources. Each entry’s key acts as a variable name that identifies the extracted data throughout the whole LCLStreamer data workflow. The entry’s associated dictionary defines the nature of the data source (via the mandatory type entry) and any other parameters needed to configure the initial data extraction from the psana library. The type of a data source is the name of a Python class within LCLStreamer that implements extraction.
For example, the following blocks declares variables timestamp and detector_data.
@@
The timestamp data class is of type Psana1Timestamp. Inside LCLStreamer, the Psana1Timestamp class will be called to read the associated timestamp data for each event. The detector_data class is instead of type Psana1AreaDetector. The two configuration parameters psana_name and calibration are passed to the Python class Psana1AreaDetector that defines how this type of data is retrieved.
LCLStreamer currently supports psana (Damiani and others, 2016) and psana2 (Thayer and others, 2024) as data sources, with a wide array of detectors and instrument readouts. It allows data to be serialized in customizable formats like HDF5 and can write them to files and/or send them via ZMQ or NNG sockets.